OpenStudy (anonymous):

Evaluate E{X(X-1)} for a poisson random variable with mean of Lambda , and hence, derive its variance

OpenStudy (anonymous):

Variance is given by \[\begin{align*}V(X)&=E(X^2)-\bigg(E(X)\bigg)^2\\ &=E(X^2)-\lambda^2\end{align*}\] Rewriting the given expression, you have \[\begin{align*}E\bigg(X(X-1)\bigg)&=E(X^2-X)\\ &=E(X^2)-E(X)\\ &=E(X^2)-\lambda \end{align*}\] You should be able to use the Poisson PDF from here to figure out the variance.